The time it takes me to wash the dishes is uniformly distributed between 12 minutes and 22 minutes. What is the probability that washing dishes tonight will take me between 18 and 20 minutes? Give your answer accurate to 2 places after the decimal point, if necessary.
Answer #1

Solution :

Given that,

a = 12

b = 22

P(c < x < d) = (d - c) / (b - a)

P(18< x < 20) = (20 - 18) / (22 - 12) =0.2

Probability = 0.2